OpenStudy (anonymous):

determine if the function f(x)=4/x+4 is one to one or not. Please give an explanation. thanks

OpenStudy (anonymous):

You have to check f(a) and f(b) 4/a+4 = 4/b+4 Does this imply that a=b than one to one. If it does not imply that a=b than its not 1to1.

OpenStudy (anonymous):

find the derivative of the function and if the derivative is always positive or always negative, the function is one to one. another way for finding the answer is to plot the function.